OpenStudy (anonymous):

Rewrite in simplest radical form 1 x −3 6 . Show each step of your process.

OpenStudy (anonymous):



OpenStudy (anonymous):

@ospreytriple

OpenStudy (anonymous):

First, I would reduce the exponent of x. What do you get?

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

Right on. So now you have\[\frac{ 1 }{ x ^{-\frac{ 1 }{ 2 }} }\]Is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK, I hope you will agree that\[\frac{ 1 }{ x ^{-\frac{ 1 }{ 2 }} } = \left( \frac{ 1 }{ x } \right)^{-\frac{ 1 }{ 2 }}\]You OK with that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Perfect. Now, to invert the base, you must change the sign of the exponent. For example,\[3^{-2} = \left( \frac{ 1 }{ 3 } \right)^{2}\] or\[\left( \frac{ 1 }{ 4 } \right)^{5} = 4^{-5}\]Gwt the idea? You need to do the same thing with\[\left( \frac{ 1 }{ x } \right)^{-\frac{ 1 }{ 2 }}\]What do you get?

OpenStudy (anonymous):

x to the pwer of 2?

OpenStudy (anonymous):

Don't invert the exponent, just change its sign. Invert the BASE, and change the sign on the EXPONENT. Try again?

OpenStudy (anonymous):

x to the power of 1/2?

OpenStudy (anonymous):

Perfect. Now write it in radical form.

OpenStudy (anonymous):

In general, fractional exponents are written in radical form like this:\[x ^{\frac{ a }{ b }} = \sqrt[b]{x ^{a}}\]

OpenStudy (anonymous):

\[\sqrt{x}\]

OpenStudy (anonymous):

Well done!

OpenStudy (anonymous):

thank you again!